A ball is thrown upward at time t=11 s on planet Ubuntu, and its acceleration from time t=11 s to t=19 s is given by the function \displaystyle a(t) = \frac{6}{t-10} \,\, \textrm{m/s}^2.

Use 4 time intervals of equal length to overestimate, and then underestimate, the relative velocity of the rising ball from time t = 11 s to t = 19 s.  
(Note: relative'' means relative to the initial velocity at t=11, which has not been given by the problem.)

(a) Overestimate
ANSWER:  m/s.

(b) Underestimate
ANSWER:
Answer #1

lets take the intervals, 11-13, 13-15,15-17, 17-19 s

overestimate relative velocity = 6*(19-11)/(19-10) = 5.33 m/s

underestimate relative velocity = 6*(13-11)/(13-10) = 4 m/s